# FedFMS: Exploring Federated Foundation Models for Medical Image Segmentation

### v1

@@ 2.1.1 SAM Architecture
SAM is a large data-driven image segmentation model. It constructs a dataset named SA-1B, consisting of 11 million images and 100 million masks, to drive its training. The image encoder utilizes a standard Vision Transformer (ViT) [2, 22] pre-trained by MAE. In our study, we set SAM using the ViT-B/16 and ViT-L/16 variants. ViT-B/16 represents the base-scale version of ViT, implemented with 768 convolutional kernels. ViT-L represents the large-scale version, implemented with 1024 convolutional kernels. The output of the image encoder is a 16× downsampled embedding of the input image. The mask decoder is a lightweight modified Transformer decoder block that includes bidirectional cross-attention and a dynamic mask prediction head.
@@ 2.2.1 Client-side Model Training
[…]
@@ 2.3 Dataset Preparation
Figure 2: The example of images and Ground Truth from different clients (A, B, …) across four datasets (a-d) and the comparison in the results of FedMSA, MSA, FedSAM and SAM. The bottom right corner of each image indicates the Dice for it or the whole nii.
@@ 3 Experiments
Table 1: The comparison of FedSAM, SAM, FedMSA, MSA, FedU-Net, and ablation variants (denoted as italics) on different medical image datasets in federated learning.
@@
Client A B C D E F Average
@@
Client B C D Average
@@
Client A B C D E F G Average
@@
3.1 Experimental Settings
3.2 Results
3.2.1 Overall Comparison
Since FedFMS is the first federated learning foundation models for medical image segmentation, there are no comparable SOTAs in this domain. We compare FedMSA with MSA and compare FedSAM with SAM. We use FedU-Net as a baseline. U-Net [19] is a commonly used and effective convolutional network for biomedical image segmentation. We use the pre-trained ViT-B architecture of SAM as default. FedMSA-L and SAM-L are extension experiments using ViT-L. Results are presented in Table 1 and Figure 2.
The performance of FedSAM, FedMSA and FedMSA-L is significantly higher than FedU-Net. Using pre-trained SAM is beneficial for medical image segmentation as it can mitigate unseen domain issues, attributed to its abundant background knowledge. The varying data quantities and distributions across clients result in inconsistent convergence directions among different clients. This inconsistency further leads to suboptimal performance of the globally aggregated model on the server. The foundation model demonstrates higher robustness and stability, which can alleviate the above issues. Therefore, through fine-tuning SAM, FedMSA, FedSAM and FedMSA-L can achieve advanced performance.
Further Discussion The test results of FedMSA and MSA are generally similar across various datasets, and the performance of FedSAM and SAM is also similar. The federated learning paradigm leads to slight differences in their performance. The discrepancies are slightly larger in their tests on clients F and G for Nuclei segmentation. The Nuclei dataset has the smallest dataset among all the datasets. Moreover, there are large differences among different types of cells. These factors lead to inconsistent convergence directions in federated learning.
@@ 3.2.2 Model Efficiency Analysis
We calculate the learnable parameter count, training time in GPU (taking the average training time in Fundus dataset as an example), GPU memory usage in each client, and the estimated FLOPs for both forward and backward propagation for FedMSA and FedSAM, as shown in Table 2. The number of parameters (denoted as $n$) to be trained and updated determines the model’s training speed and communication cost. The amount of parameters to be communicated in each round of federated learning is $2n$. The results show that FedMSA freezes a large number of parameters in the encoder, resulting in a significantly reduced parameter count and FLOPs compared to FedSAM, and consequently reducing communication and training costs.
@@
Index Learnable Parameter Average Training Time GPU Memory Usage FLOPs
Model FedMSA FedSAM FedMSA FedSAM FedMSA FedSAM FedMSA FedSAM
Value 14.7 B 93.7 B 739.9 min 911.4 min 52,274 MiB 58,478 MiB 5.7 T 13.4 T
@@ 3.2.3 Pre-training Impact
SAM is pre-trained on a large-scale natural dataset. To investigate the effectiveness of this pretraining for medical image segmentation and its impact on federated learning, we conducted an ablation study. For FedSAM, FedMSA, SAM and MSA, we constructed variants FedSAM (-PT), FedMSA (-PT), SAM (-PT) and MSA (-PT) without using pre-trained parameters, and the experimental results are illustrated in Table 1. The results show that not using pre-trained parameters from SAM leads to a drastic decrease in performance. In some cases, FedSAM (-PT) performs worse than SAM (-PT), which is due to the convergence of no pre-trained SAM under federated learning is not stable. FedSAM (-PT) and FedMSA (-PT) sometimes perform worse than the lightweight model FedU-Net, for example in experiments A, B, C, D, F, G on Nuclei dataset and B, D, E, F on Prostate Cancer dataset. This indicates that the pretraining knowledge of SAM is crucial for its effectiveness under the federated learning paradigm, which enables our federated foundation models to far surpass the traditional federated learning model (FedU-Net).

### v2

@@ Abstract
Keywords:
Medical image segmentation Federated learning Foundation model.
@@ 2.1.1 SAM Architecture
SAM is a large data-driven image segmentation model. It constructs a dataset named SA-1B, consisting of 11 million images and 100 million masks, to drive its training. The image encoder utilizes a standard Vision Transformer (ViT) [2, 23] pre-trained by Masked Autoencoders. In our study, we set SAM using the ViT-B/16 and ViT-L/16 variants. ViT-B/16 represents the base-scale version of ViT, implemented with 768 convolutional kernels. ViT-L represents the large-scale version, implemented with 1024 convolutional kernels. The output of the image encoder is a 16× downsampled embedding of the input image. The mask decoder is a lightweight modified Transformer decoder block that includes bidirectional cross-attention and a dynamic mask prediction head.
@@ 2.2.1 Client-side Model Training
FedMSA For MSA, we fine-tune the parameters of adapters in the encoder (denoted as $W^{(a)}_{t}$), and all parameters in the decoder. Our MSA’s decoder adopts the same multi-class segmentation decoder as our SAM. The features obtained by fine-tuning the encoder propagate to the top layers of the decoder, so all parameters of the decoder need to be fine-tuned. The parameter of the SAM decoder is lightweight, resulting in a low fine-tuning cost. The structure of MSA is shown in Figure 1 (b). We use MSA to build a more efficient federated learning framework for three reasons. (1) MSA performs well in fine-tuning tasks for medical image segmentation. (2) Only training adapters and decoder requires less computational cost compared to training the entire SAM. (3) During global parameter aggregation, only the parameters of the adapters and decoder need to be transmitted and calculated. During the federated $t$-th round, the local model is initialized by fetching global model parameters $W_{t}=\{W^{(d)}_{t},W^{(a)}_{t}\}$ from the server. In the $k$-th client, the updated parameters is $W_{t+1}^{k}=\{W^{(d)}_{t+1,k},W^{(a)}_{t+1,k}\}$, where $W^{(a)}_{t+1,k}$ is the updated parameters of the adapters.
@@ 2.3 Dataset Preparation
Figure 2: The example of images and Ground Truth from different clients across four datasets (a-d) and the comparison in the results of FedMSA, MSA, FedSAM and SAM. The bottom right corner of each image indicates the Dice for it or the whole nii.
@@ 3 Experiments
3.1 Experimental Settings
3.2 Results
3.2.1 Overall Comparison
To explore the impact of federated learning on foundation model, we compare FedMSA with MSA and compare FedSAM with SAM. We use FedU-Net and FednnU-Net as baselines. U-Net [20] is a commonly used and effective convolutional network for biomedical image segmentation. nnU-Net [6] is a more robust method compared to U-Net. We use the pre-trained ViT-B architecture of SAM as default. FedMSA-L and SAM-L are extension experiments using ViT-L. Results are presented in Table 1 and Figure 2.
The performance of FedSAM, FedMSA and FedMSA-L is significantly higher than FedU-Net and FednnU-Net. The preprocessing of nnU-Net under federated learning is limited in effectiveness, and its training on Non-IID datasets is also unstable. Using pre-trained SAM is beneficial for medical image segmentation as it can mitigate unseen domain issues, attributed to its abundant background knowledge. The varying data quantities and distributions across clients result in inconsistent convergence directions among different clients. This inconsistency further leads to suboptimal performance of the globally aggregated model on the server. The foundation model demonstrates higher robustness and stability, which can alleviate the above issues. Therefore, through fine-tuning SAM, FedMSA, FedSAM and FedMSA-L can achieve advanced performance.
Table 1: The comparison of FedSAM, SAM, FedMSA, MSA, FedU-Net, and ablation variants (denoted as italics) on different medical image datasets. The p-values between FedSAM and SAM, as well as between FedMSA and MSA, are both greater than 0.5.
@@ 3.2.1 Overall Comparison
Client BIDMC HK I2CVB ISBI ISBI1.5 UCL Average
@@
Client ORIGA G1020 Drishit-GS1 Average
@@
Client Adrenal Esophagus Bile-duct Uterus MoNuSAC TNBC MoNuSeg Average
@@
Further Discussion The test results of FedMSA and MSA are generally similar across various datasets, and the performance of FedSAM and SAM is also similar. The federated learning paradigm leads to slight differences in their performance. The discrepancies are slightly larger in their tests on clients TNBC and MoNuSeg for Nuclei segmentation. The Nuclei dataset has the smallest dataset among all the datasets. Moreover, there are large differences among different types of cells. These factors lead to inconsistent convergence directions in federated learning.
@@ 3.2.2 Model Efficiency Analysis
We calculate the learnable parameter count, training time in GPU (the average training time in Fundus dataset), GPU memory usage in each client, and the estimated FLOPs (Floating Point Operations) for both forward and backward propagation for FedMSA and FedSAM, as shown in Table 2. The number of parameters (denoted as $n$) to be trained and updated determines the model’s training speed and communication cost. The amount of parameters to be communicated in each round of federated learning is $2n$. The results show that FedMSA freezes a large number of parameters in the encoder, resulting in a significantly reduced parameter count and FLOPs compared to FedSAM, and consequently reducing communication and training costs. We calculate the average time required to predict each 2D image, as shown in "Predicting time" in Table 2. During the prediction process, FedMSA takes a bit longer because it has more parameters from adapters compared to FedSAM.
@@
Learnable Parameter Training Time GPU Memory Usage FLOPs Predicting Time
FedMSA FedSAM FedMSA FedSAM FedMSA FedSAM FedMSA FedSAM FedMSA FedSAM
14.7 B 93.7 B 739.9 min 911.4 min 52,274 MiB 58,478 MiB 5.7 T 13.4 T 0.127 s 0.118 s
@@ 3.2.3 Pre-training Impact
SAM is pre-trained on a large-scale natural dataset. To investigate the effectiveness of this pretraining for medical image segmentation and its impact on federated learning, we conducted an ablation study. For FedSAM, FedMSA, SAM and MSA, we constructed variants FedSAM (-PT), FedMSA (-PT), SAM (-PT) and MSA (-PT) without using pre-trained parameters, and the experimental results are illustrated in Table 1. The results show that not using pre-trained parameters from SAM leads to a drastic decrease in performance. In some cases, FedSAM (-PT) performs worse than SAM (-PT), which is due to the convergence of no pre-trained SAM under federated learning is not stable. FedSAM (-PT) and FedMSA (-PT) sometimes perform worse than the lightweight model FedU-Net, for example in experiments Adrenal, Esophagus, Bile-duct, Uterus, TNBC, MoNuSeg on Nuclei dataset and HK, ISBI, ISBI1.5, UCL on Prostate Cancer dataset. This indicates that the pretraining knowledge of SAM is crucial for its effectiveness under the federated learning paradigm, which enables our federated foundation models to far surpass the traditional federated learning models (e.g., FednnU-Net). The code for our implementation of FednnU-Net is available at https://github.com/LMIAPC/FednnU-Net.
@@ 4 Conclusion
{credits}
4.0.1 Acknowledgements
This work is supported by Shenzhen Science and Technology Program (No.JCYJ20230807120800001), and 2023 Shenzhen sustainable supporting funds for colleges and universities (No.20231121165240001). The authors sincerely appreciate the computing environment supported by the China Unicom Shenzhen Intelligent Computing Center.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.